OpenStudy (anonymous):

Solve the equation: x^4-x^3-x^2-x-2=0

OpenStudy (anonymous):

You may try to spot roots by eye or by the rational roots theorem to begin...

OpenStudy (anonymous):

yes, but it isn't easy :(

OpenStudy (anonymous):

Do you know rational roots theorem?

OpenStudy (anonymous):

unfortunately not :(

OpenStudy (anonymous):

OK, if you have a polynomial ax^n+....+c then possible rational roots are given by plus/minus factor combinations of c/a In this case, 2/1 so 1/1, -1/1, 2/2, -2/2 so you shoudl first check plus minus 1 and 2 (actually it is a good idea to check plus minus 1 anyway)

hartnn (hartnn):

may i suggest a way to find one of the roots easily: The rule below will help in determining whether x=1 and/or x=-1 is a root of the equation or not: 1.Find the sum of coefficients of odd power terms(coefficients of x^3 and x here) 2.Find the sum of coefficients of even power terms(coefficient of x^4,x^2 and constant) If these two are equal then (x=−1) is the root and (x+1) is the factor. If the sum of these two=0, then (x=1) is the root and (x−1) is the factor. Once we get x=+1 and/or x=-1 as roots we can always do synthetic division or divide the given polynomial by (x±1)

OpenStudy (anonymous):

so how it should be write ? x^4-1-x^3+1-x^2+1-x+1?

OpenStudy (anonymous):

x^4-x^3-x^2-x-2=0 (-1)^4 -(-1)^3 -(-1)^2 - (-1) -2 = 0 is true so x = -1 works If you test 2, you will find that works as well. So then you have (x+1)(x-2)(quadratic) = 0 left to solve which is not too hard.

hartnn (hartnn):

to get the quadratic, u need to do synthetic division or long division of polynomials.

hartnn (hartnn):

by dividing given polynomial by x+1 and then x-2

OpenStudy (anonymous):

but how did you this (x+1)(x-2)(quadratic) = 0 ?

OpenStudy (anonymous):

I wrote it above??

OpenStudy (anonymous):

ok, but how i should to know that 1 is good ?

OpenStudy (anonymous):

sorry -1 ?

OpenStudy (anonymous):

What I wrote: "(-1)^4 -(-1)^3 -(-1)^2 - (-1) -2 = 0 is true so x = -1 works"

OpenStudy (anonymous):

so i must count for 0, 1, -1 and looking for a good number?

OpenStudy (anonymous):

Again, what I wrote: "OK, if you have a polynomial ax^n+....+c then possible rational roots are given by plus/minus factor combinations of c/a In this case, 2/1 so 1/1, -1/1, 2/2, -2/2 so you shoudl first check plus minus 1 and 2 (actually it is a good idea to check plus minus 1 anyway)"

OpenStudy (anonymous):

ok,now i understand :) thank you very very much :)

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

:)